OpenStudy (anonymous):

A $200.00 sweater has been marked down to $120.00. What percentage discount from the original price does the markdown represent?

OpenStudy (anonymous):

I always set up proportions to solve problems for percents, have you learned that way?

OpenStudy (anonymous):

\[80/200=x/100\] \[.4=x/100\] \[x=40%\]

OpenStudy (anonymous):

In the numerator you put the part of the whole that you're looking at (in this case we took $80 off the price of the sweater). In the denominator you put the entire amount ($200). On the other side of the equals sign you put x/100, because percentages are parts out of 100. Then solve for x.